Let W {(x, y, z) R3: x + y + z = 0}. Check
if W is a subspace of R3 . Find a non-zero
subspace U of R3 so that W(intersection)U = (0).
1
Expert's answer
2020-03-23T15:50:22-0400

"W=\{(x,y,z)\in\mathbb{R}^3: x+y+z=0\}"

1) "u=(0,0,0): \ u\in W?"

"0+0+0=0\ \Rightarrow u\in W"

2) "u=(x_1,y_1,z_1), \ v=(x_2,y_2,z_2), u,v\in W: \ \ u+v\in W?"

"u+v=(x_1+x_2,y_1+y_2,z_1+z_2)"

"(x_1+x_2)+(y_1+y_2)+(z_1+z_2)=(x_1+y_1+z_1)+(x_2+y_2+z_2)=0+0=0"

"\Rightarrow \ u+v\in W"

3) "u=(x,y,z)\in W, \lambda \in \mathbb{R}: \ \lambda u\in W?"

"\lambda x+\lambda y+\lambda z=\lambda(x+y+z)=\lambda \cdot 0=0\Rightarrow \ \lambda u\in W"

Therefore, "W" is a subspace of "\mathbb{R}^3"


"U=\{(x,y,z)\in \mathbb{R}^3: x=y=z\}" , "U" is non-zero because "(1,1,1)\in U"

1) "u=(0,0,0): \ u\in U?"

"0=0=0 \ \Rightarrow u\in U"

2) "u=(x_1,y_1,z_1), \ v=(x_2,y_2,z_2), u,v\in U: \ \ u+v\in U?"

"x_1=y_1=z_1, \ x_2=y_2=z_2\ \Rightarrow x_1+x_2=y_1+y_2=z_1+z_2\ \Rightarrow u+v\in U"

3) "u=(x,y,z)\in U, \lambda \in \mathbb{R}: \ \lambda u\in U?"

"x=y=z\ \Rightarrow \lambda x
=\lambda y=\lambda z\ \Rightarrow \lambda u\in U"

Therefore, "U" is a (non-zero) subspace of "\mathbb{R}^3"


Suppose that "u\in W \cap U"

"\begin{cases}
x+y+z=0
\\
x=y=z
\end{cases}
\quad
\begin{cases}
3x=0
\\
x=y=z
\end{cases} \Rightarrow x=y=z=0, \ \ u=(0,0,0)"

Therefore, "W\cap U=\{0\}"